# Residual-based Chebyshev filtered subspace iteration for sparse Hermitian eigenvalue problems tolerant to inexact matrix-vector products

### v3

@@ 3.2.1 Traditional Chebyshev filtering method employing low-precision matrix-products
If $\textbf{H}\in\mathbb{C}^{m\times m}$ and $\textbf{X}\in\mathbb{C}^{m\times n}$, the spectral norm of the error in evaluating the matrix product HX due to floating point approximations satisfies $-{\textbf{H}\otimes\textbf{X}}\rVert\,\leq\gamma_{m}\lVert\textbf{H}\rVert\,\lVert\textbf{X}\rVert\,$ H ∥ ∥ X ∥
Proof 3.4.
@@
This is a straightforward application of the results derived in [14, section 3.5]. We note that $\gamma_{m}$ is a constant that depends on $m$, and the machine precision of the floating point arithmetic (denoted as $\varepsilon_{M}$). For the case of dense real matrices, we have $\gamma_{m}=m\varepsilon_{M}/(1-m\varepsilon_{M})$ and for dense complex matrices, it needs to be appropriately modified as described in [14, section 3.6]. Further, we also note that in the case where H is sparse, significantly tighter bounds independent of $m$ can be achieved depending on the specific implementations of the sparse matrix-dense matrix multiplication routines.
Lemma 3.5.
@@
(8) $\boldsymbol{\Delta}_{k+1}^{(i)}=a_{k}\textbf{H}\boldsymbol{\Delta}_{k}^{(i)}+b_{k}\boldsymbol{\Delta}_{k}^{(i)}+c_{k}\boldsymbol{\Delta}_{k-1}^{(i)}+a_{k}\textbf{e}_{k}^{(i)}$
@@
(9) $\lVert\boldsymbol{\Delta}_{k+1}^{(i)}\rVert\,\leq g_{k}^{0}\left(\lVert\boldsymbol{\Delta}_{0}^{(i)}\rVert\,+\lVert\boldsymbol{\Delta}_{1}^{(i)}\rVert\,\right)+g_{k}^{1}h_{1}^{(i)}$
Proof 3.6.
@@
We define the following
$\textbf{d}_{k}^{(i)}$ $=\begin{bmatrix}\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,\\ \lVert\boldsymbol{\Delta}_{k-1}^{(i)}\rVert\,\end{bmatrix}$ $\textbf{F}_{k}$ $=\begin{bmatrix}\lvert a_{k}\rvert\,\lVert\textbf{H}\rVert\,+\lvert b_{k}\rvert\,+\lvert a_{k}\rvert\,h_{0}&\lvert c_{k}\rvert\,\\ 1&0\end{bmatrix}$ $\textbf{E}_{k}^{(i)}$ $=\begin{bmatrix}\lvert a_{k}\rvert\,h_{1}^{(i)}\\ 0\end{bmatrix}$
@@
Defining $g_{k}^{0}=\prod_{j=1}^{k}\lVert\textbf{F}_{j}\rVert\,$ and $g_{k}^{1}=\sum_{j=1}^{k}\prod_{r=j+1}^{k}\lVert\textbf{F}_{r}\rVert\,\lvert a_{j}\rvert\,$ we have
$\lVert\boldsymbol{\Delta}_{k+1}^{(i)}\rVert\,\leq g_{k}^{0}\left(\lVert\boldsymbol{\Delta}_{0}^{(i)}\rVert\,+\lVert\boldsymbol{\Delta}_{1}^{(i)}\rVert\,\right)+g_{k}^{1}h_{1}^{(i)}$
Theorem 3.7.
The spectral norm of the error $\boldsymbol{\Delta}_{k}^{(i)}=\underline{\textbf{Y}}_{k}^{(i)}-\textbf{Y}_{k}^{(i)}$ in the subspace construction using lower precision arithmetic in matrix-products of the recurrence relation Eq. 2 satisfies $\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,\leq\gamma_{m}\eta_{k}$ for $k=0,1,\dots,p$ where $\eta_{k}$ is some finite constant that depends on $k$
Proof 3.8.
Using Eq. 2 and Eq. 7 we can now write a recurrence relation for $\boldsymbol{\Delta}_{k}^{(i)}=\underline{\textbf{Y}}_{k}^{(i)}-\textbf{Y}_{k}^{(i)}$ as
(10) $\boldsymbol{\Delta}_{k+1}^{(i)}=a_{k}\textbf{H}\boldsymbol{\Delta}_{k}^{(i)}+b_{k}\boldsymbol{\Delta}_{k}^{(i)}+c_{k}\boldsymbol{\Delta}_{k-1}^{(i)}+a_{k}\left({\textbf{H}\otimes\underline{\textbf{Y}}_{k}^{(i)}}-\textbf{H}\underline{\textbf{Y}}_{k}^{(i)}\right)$
with the initial conditions $\boldsymbol{\Delta}_{0}^{(i)}=\boldsymbol{\Delta}_{1}^{(i)}=0$. We note that this recurrence relation is of the same form as that of Eq. 8 in Lemma 3.5 with $\textbf{e}_{k}^{(i)}={\textbf{H}\otimes\underline{\textbf{Y}}_{k}^{(i)}}-\textbf{H}\underline{\textbf{Y}}_{k}^{(i)}$. From Lemma 3.3 we have $\lVert\textbf{e}_{k}^{(i)}\rVert\,=\lVert{\textbf{H}\otimes\underline{\textbf{Y}}_{k}^{(i)}}-\textbf{H}\underline{\textbf{Y}}_{k}^{(i)}\rVert\,\leq\gamma_{m}\lVert\textbf{H}\rVert\,\lVert\underline{\textbf{Y}}_{k}^{(i)}\rVert\,$. We further note that
$\lVert\underline{\textbf{Y}}_{k}^{(i)}\rVert\,=\lVert\boldsymbol{\Delta}_{k}^{(i)}+\textbf{Y}_{k}^{(i)}\rVert\,\leq\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,+\lVert\textbf{Y}_{k}^{(i)}\rVert\,\leq\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,+\lVert C_{k}(\textbf{H})\textbf{X}^{(i)}\rVert\,\\ \leq\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,+\lVert C_{k}(\textbf{H})\rVert\,$
Thus we have $\lVert\textbf{e}_{k}^{(i)}\rVert\,\leq\gamma_{m}\lVert\textbf{H}\rVert\,\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,+\gamma_{m}\lVert\textbf{H}\rVert\,\lVert C_{k}(\textbf{H})\rVert\,$. Using Lemma 3.5 with $\lVert\boldsymbol{\Delta}_{0}^{(i)}\rVert\,=\lVert\boldsymbol{\Delta}_{1}^{(i)}\rVert\,=0$, $h_{0}=\gamma_{m}\lVert\textbf{H}\rVert\,$ and $h_{1}^{(i)}=\gamma_{m}\lVert\textbf{H}\rVert\,\lVert C_{k}(\textbf{H})\rVert\,$ we obtain $\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,\leq\gamma_{m}\eta_{k}$ with $\eta_{k}=\lVert\textbf{H}\rVert\,\lVert C_{k}(\textbf{H})\rVert\,g_{k-1}^{1}$.
(11) $\lvert C_{p}(\lambda_{n})\rvert\,-\lvert C_{p}(\lambda_{n+1})\rvert\,$ $>\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,\left(\sec{\angle(\mathcal{S}^{(i)},\mathcal{S})}+\csc{\angle(\mathcal{S}^{(i)},\mathcal{S})}\right)$
While the above inequality is satisfied, $\angle(\mathcal{S}^{(i)},\mathcal{S})$ reduces as $i$ increases (from Theorem 3.1). However, we note that the RHS of this equation does not monotonically decrease with a decrease in $\angle(\mathcal{S}^{(i)},\mathcal{S})$ and increases as $\angle(\mathcal{S}^{(i)},\mathcal{S})$ approaches 0. Thus, it stands to reason that beyond a certain value of $\angle(\mathcal{S}^{(i)},\mathcal{S})$, the inequality in Eq. 11 no longer holds, and we cannot approach the wanted eigenspace beyond this point. Under the condition that $\gamma_{m}\eta_{p}<<\lvert C_{p}(\lambda_{n})\rvert\,-\lvert C_{p}(\lambda_{n+1})\rvert\,$, we can estimate the closest angle that can be achieved and is given by the following expression
(12) $\angle(\mathcal{S}^{(i)},\mathcal{S})\approx\frac{\gamma_{m}\eta_{p}}{\lvert C_{p}(\lambda_{n})\rvert\,-\lvert C_{p}(\lambda_{n+1})\rvert\,}$
3.2.2 Proposed residual-based Chebyshev filtering approach
We now propose a residual-based reformulation of the recurrence relation defined in Eq. 2. To this end we define the residual $\textbf{R}^{i}_{k}$ in a given iteration $i$ for $k=0,\dots,p$ in the following way:
(13) $\textbf{R}_{k}^{(i)}=C_{k}(\textbf{H})\textbf{X}^{(i)}-\textbf{X}^{(i)}C_{k}(\boldsymbol{\Lambda}^{(i)})=\textbf{Y}_{k}^{(i)}-\textbf{X}^{(i)}\boldsymbol{\Lambda}_{k}^{(i)}\;\;\;\text{for}\;\;\;k=0,\dots,p$
where we have defined $\boldsymbol{\Lambda}_{k}^{(i)}=C_{k}(\boldsymbol{\Lambda}^{(i)})$ and recall $p$ is the maximum Chebyshev polynomial degree used in the subspace construction step.
Proposition 3.9.
The recurrence relation given by Eq. 2 can be reformulated in terms of the residuals defined by $\textbf{R}_{k}^{(i)}=\textbf{Y}_{k}^{(i)}-\textbf{X}^{(i)}\boldsymbol{\Lambda}_{k}^{(i)}$ as
(14) $\textbf{R}_{k+1}^{(i)}$ $=a_{k}\textbf{H}\textbf{R}_{k}^{(i)}+b_{k}\textbf{R}_{k}^{(i)}+c_{k}\textbf{R}_{k-1}^{(i)}+a_{k}\textbf{R}^{(i)}\boldsymbol{\Lambda}_{k}^{(i)}$
(15) $\boldsymbol{\Lambda}_{k+1}^{(i)}$ $=a_{k}\boldsymbol{\Lambda}_{k}^{(i)}\boldsymbol{\Lambda}^{(i)}+b_{k}\boldsymbol{\Lambda}_{k}^{(i)}+c_{k}\boldsymbol{\Lambda}_{k-1}^{(i)}$
where $\textbf{R}_{0}^{(i)}=0$ and $\textbf{R}_{1}^{(i)}=\frac{\sigma_{1}}{e}\textbf{R}^{(i)}$ with $\textbf{R}^{(i)}=\textbf{H}\textbf{X}^{(i)}-\textbf{X}^{(i)}\boldsymbol{\Lambda}^{(i)}$ and further we have $a_{k}=2\sigma_{k+1}/e$, $b_{k}=-2\sigma_{k+1}c/e$, and $c_{k}=-\sigma_{k}\sigma_{k+1}$
Proof 3.10.
Multiplying with $\textbf{X}^{(i)}$ and subtracting from Eq. 2, we have
$\textbf{R}_{k+1}^{(i)}$ $=a_{k}\textbf{H}\textbf{R}_{k}^{(i)}+b_{k}\textbf{R}_{k}^{(i)}+c_{k}\textbf{R}_{k-1}^{(i)}+a_{k}\textbf{R}^{(i)}\boldsymbol{\Lambda}_{k}^{(i)}$
After computing $\textbf{R}_{p}^{(i)}$ using this recurrence relation we can now recover the desired subspace $\textbf{Y}_{p}^{(i)}$ using the relation in Eq. 13 i.e., $\textbf{Y}_{p}^{(i)}=\textbf{R}_{p}^{(i)}+\textbf{X}^{(i)}\boldsymbol{\Lambda}_{p}^{(i)}$. We note that using approximate matrix multiplication for evaluating $\textbf{H}\textbf{R}_{k}^{(i)}$ we can write the following recurrence relation
(16) $\underline{\textbf{R}}_{k+1}^{(i)}=a_{k}{\textbf{H}\otimes\underline{\textbf{R}}_{k}^{(i)}}+b_{k}\underline{\textbf{R}}_{k}^{(i)}+c_{k}\underline{\textbf{R}}_{k-1}^{(i)}+a_{k}\textbf{R}^{(i)}\boldsymbol{\Lambda}_{k}^{(i)}$
and consequently, we have $\underline{\textbf{Y}}_{p}^{(i)}=\underline{\textbf{R}}_{p}^{(i)}+\textbf{X}^{(i)}\boldsymbol{\Lambda}_{p}^{(i)}$.
We now prove a result useful for further analysis
Lemma 3.11.
@@ 3.2.2 Proposed residual-based Chebyshev filtering approach
where we have defined $f_{k}=\sum_{j=0}^{k}\sum_{r=0}^{j-1}\lVert\alpha_{j}\textbf{H}^{j-r-1}\rVert\,\lVert{\boldsymbol{\Lambda}^{(i)}}^{r}\rVert\,$.
Lemma 3.13.
The spectral norm of the error $\boldsymbol{\Delta}_{k}^{(i)}=\underline{\textbf{Y}}_{k}^{(i)}-\textbf{Y}_{k}^{(i)}$ in the subspace construction using lower precision arithmetic in matrix-products of the recurrence relation Eq. 16 satisfies $\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,\leq\gamma_{m}\tilde{\eta}_{k}\lVert\textbf{R}^{(i)}\rVert\,$ for $k=0,1,\dots,p$ where $\tilde{\eta}_{k}$ is some finite constant that depends on $k$.
Proof 3.14.
Using Eq. 14 and Eq. 16 we can now write a recurrence relation for $\boldsymbol{\Delta}_{k}^{(i)}=\underline{\textbf{Y}}_{k}^{(i)}-\textbf{Y}_{k}^{(i)}$ as
(17) $\boldsymbol{\Delta}_{k+1}^{(i)}=a_{k}\textbf{H}\boldsymbol{\Delta}_{k}^{(i)}+b_{k}\boldsymbol{\Delta}_{k}^{(i)}+c_{k}\boldsymbol{\Delta}_{k-1}^{(i)}+a_{k}\left({\textbf{H}\otimes\underline{\textbf{R}}_{k}^{(i)}}-\textbf{H}\underline{\textbf{R}}_{k}^{(i)}\right)$
with the initial conditions $\boldsymbol{\Delta}_{0}^{(i)}=\boldsymbol{\Delta}_{1}^{(i)}=0$. We note that this recurrence relation is of the same form as that of Eq. 8 in Lemma 3.5 with $\textbf{e}_{k}^{(i)}={\textbf{H}\otimes\underline{\textbf{R}}_{k}^{(i)}}-\textbf{H}\underline{\textbf{R}}_{k}^{(i)}$. From Lemma 3.3 we have $\lVert\textbf{e}_{k}^{(i)}\rVert\,=\lVert{\textbf{H}\otimes\underline{\textbf{R}}_{k}^{(i)}}-\textbf{H}\underline{\textbf{R}}_{k}^{(i)}\rVert\,\leq\gamma_{m}\lVert\textbf{H}\rVert\,\lVert\underline{\textbf{R}}_{k}^{(i)}\rVert\,$. We further note that
$\lVert\underline{\textbf{R}}_{k}^{(i)}\rVert\,=\lVert\boldsymbol{\Delta}_{k}^{(i)}+\textbf{R}_{k}^{(i)}\rVert\,\leq\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,+\lVert\textbf{R}_{k}^{(i)}\rVert\,\leq\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,+f_{k}\lVert\textbf{R}^{(i)}\rVert\,$
Thus we have $\lVert\textbf{e}_{k}^{(i)}\rVert\,\leq\gamma_{m}\lVert\textbf{H}\rVert\,\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,+\gamma_{m}f_{k}\lVert\textbf{H}\rVert\,\lVert\textbf{R}^{(i)}\rVert\,$. Using Lemma 3.5 with $\lVert\boldsymbol{\Delta}_{0}^{(i)}\rVert\,=\lVert\boldsymbol{\Delta}_{1}^{(i)}\rVert\,=0$, $h_{0}=\gamma_{m}\lVert\textbf{H}\rVert\,$ and $h_{1}^{(i)}=\gamma_{m}f_{k}\lVert\textbf{H}\rVert\,\lVert\textbf{R}^{(i)}\rVert\,$ we obtain $\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,\leq\gamma_{m}\tilde{\eta}_{k}\lVert\textbf{R}^{(i)}\rVert\,$ with $\tilde{\eta}_{k}=f_{k}\lVert\textbf{H}\rVert\,g_{k-1}^{1}$.
Lemma 3.15.
@@
(18) $\lVert\textbf{R}^{(i)}\rVert\,$ $\leq\sin{\angle(\mathcal{S}^{(i)},\mathcal{S})}(\lVert\boldsymbol{\Lambda}_{1}\rVert\,+\lVert\boldsymbol{\Lambda}_{2}\rVert\,)\leq 2\lVert\textbf{H}\rVert\,\sin{\angle(\mathcal{S}^{(i)},\mathcal{S})}$
Theorem 3.17.
(19) $\lvert C_{p}(\lambda_{n})\rvert\,-\lvert C_{p}(\lambda_{n+1})\rvert\,$ $>2\lVert\textbf{H}\rVert\,\gamma_{m}\tilde{\eta}_{p}\left(1+\tan{\angle(\mathcal{S}^{(i)},\mathcal{S})}\right)$
Proof 3.18.
From Lemmas 3.11, 3.13, and 3.15 we have $\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,\leq 2\lVert\textbf{H}\rVert\,\gamma_{m}\tilde{\eta}_{p}\sin{\angle(\mathcal{S}^{(i)},\mathcal{S})}$ and in order to have $\angle(\underline{\mathcal{S}}^{(i+1)},\mathcal{S})<\angle(\mathcal{S}^{(i)},\mathcal{S})$ from Eq. 3 we require
(20) $\lvert C_{p}(\lambda_{n})\rvert\,-\lvert C_{p}(\lambda_{n+1})\rvert\,$ $>2\lVert\textbf{H}\rVert\,\gamma_{m}\tilde{\eta}_{p}\left(1+\tan{\angle(\mathcal{S}^{(i)},\mathcal{S})}\right)$
We note that if this inequality is satisfied for some $i=i_{0}$ then it holds for all $i>i_{0}$ as the RHS decreases with decreasing $\angle(\mathcal{S}^{(i)},\mathcal{S})$ and hence we conclude that the R-ChFSI method can converge under approximations where the ChFSI method fails.
3.3 Generalized Eigenvalue Problems
We now analyze the case of a Hermitian generalized eigenproblem in Eq. 1 given by $\textbf{A}\textbf{u}_{i}=\lambda_{i}\textbf{B}\textbf{u}_{i}$ or equivalently $\textbf{H}\textbf{u}_{i}=\lambda_{i}\textbf{u}_{i}$ with $\textbf{H}=\textbf{B}^{-1}\textbf{A}$. In constructing a subspace rich in the desired eigenvectors using Chebyshev filtered subspace iteration approach, we consider the specific case of approximating $\textbf{B}^{-1}$ with a matrix $\textbf{D}^{-1}$ such that $\textbf{D}^{-1}\textbf{B}\approx\textbf{I}$ in evaluating the matrix product $\textbf{H}\textbf{Y}^{(i)}_{k}$ within the Chebyshev recurrence relation defined by Eq. 2. We note that specific forms of the approximation, such as choosing D to be diagonal or block diagonal, can give significant performance benefits.
3.3.1 Traditional Chebyshev filtering approach with inexact matrix-products involving $\textbf{D}^{-1}$
Using the approximation $\textbf{D}^{-1}$ for $\textbf{B}^{-1}$, the construction of filtered subspace using the recurrence relation in Eq. 2 is now accomplished by defining $\underline{\textbf{Y}}^{(i)}_{k}={C_{k}(\textbf{D}^{-1}\textbf{A})\textbf{X}^{(i)}}$, where $k=2,\dots,p$, and subsequently the 2-term recurrence relation can be written as
(21) $\underline{\textbf{Y}}_{k+1}^{(i)}=a_{k}{\textbf{D}^{-1}\textbf{A}\underline{\textbf{Y}}_{k}^{(i)}}+b_{k}\underline{\textbf{Y}}_{k}^{(i)}+c_{k}\underline{\textbf{Y}}_{k-1}^{(i)}$
with the initial conditions $\underline{\textbf{Y}}_{0}^{(i)}=\textbf{X}^{(i)}$ and $\underline{\textbf{Y}}_{1}^{(i)}=\sigma_{1}\frac{\textbf{D}^{-1}\textbf{A}-c\textbf{I}}{e}\textbf{X}^{(i)}$.
Theorem 3.19.
The spectral norm of the error $\boldsymbol{\Delta}_{k}^{(i)}=\underline{\textbf{Y}}_{k}^{(i)}-\textbf{Y}_{k}^{(i)}$ in the subspace construction using inexact matrix products of the recurrence relation Eq. 2 involving $\textbf{D}^{-1}$ satisfies $\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,\leq\zeta\eta_{k}$ for $k=0,1,\dots,p$ where $\eta_{k}$ is some finite constant that depends on $k$ and $\zeta=\lVert\textbf{D}^{-1}-\textbf{B}^{-1}\rVert\,$
Proof 3.20.
@@ 3.3.1 Traditional Chebyshev filtering approach with inexact matrix-products involving $\textbf{D}^{-1}$
Using Eq. 2 and Eq. 21 we can now write a recurrence relation for $\boldsymbol{\Delta}_{k}^{(i)}=\underline{\textbf{Y}}_{k}^{(i)}-\textbf{Y}_{k}^{(i)}$ as
(22) $\boldsymbol{\Delta}_{k+1}^{(i)}=a_{k}\textbf{H}\boldsymbol{\Delta}_{k}^{(i)}+b_{k}\boldsymbol{\Delta}_{k}^{(i)}+c_{k}\boldsymbol{\Delta}_{k-1}^{(i)}+a_{k}\left({\textbf{D}^{-1}\textbf{A}\underline{\textbf{Y}}_{k}^{(i)}}-\textbf{B}^{-1}\textbf{A}\underline{\textbf{Y}}_{k}^{(i)}\right)$

### v4

@@ A Useful Lemmas
If $\textbf{H}\in\mathbb{C}^{m\times m}$ and $\textbf{X}\in\mathbb{C}^{m\times n}$, the spectral norm of the error in evaluating the matrix product HX due to floating point approximations satisfies $\lVert\textbf{H}\textbf{X}-{\textbf{H}\otimes\textbf{X}}\rVert\,\leq\gamma_{m}\lVert\textbf{H}\rVert\,\lVert\textbf{X}\rVert\,$
Proof.
@@
This is a straightforward application of the results derived in [highamAccuracyStabilityNumerical2002, section 3.5]. We note that $\gamma_{m}$ is a constant that depends on $m$, and the machine precision of the floating point arithmetic (denoted as $\varepsilon_{M}$). For the case of dense real matrices, we have $\gamma_{m}=m^{2}\varepsilon_{M}/(1-m\varepsilon_{M})$ and for dense complex matrices, it needs to be appropriately modified as described in [highamAccuracyStabilityNumerical2002, section 3.6]. We also note that tigher bounds can be obtained from probabilistic error analysis [ipsenProbabilisticErrorAnalysis2020, highamSharperProbabilisticBackward2020]. Further, we also note that in the case where H is sparse, significantly tighter bounds independent of $m$ can be achieved depending on the specific implementations of the sparse matrix-dense matrix multiplication routines. ∎
Lemma A.3.
@@
$\boldsymbol{\Delta}_{k+1}^{(i)}=a_{k}\textbf{H}\boldsymbol{\Delta}_{k}^{(i)}+b_{k}\boldsymbol{\Delta}_{k}^{(i)}+c_{k}\boldsymbol{\Delta}_{k-1}^{(i)}+a_{k}\textbf{e}_{k}^{(i)}$ (16)
@@
$\lVert\boldsymbol{\Delta}_{k+1}^{(i)}\rVert\,\leq g_{k}^{0}\left(\lVert\boldsymbol{\Delta}_{0}^{(i)}\rVert\,+\lVert\boldsymbol{\Delta}_{1}^{(i)}\rVert\,+g_{k}^{1}h_{1}^{(i)}\right)$ (17)
Proof.
@@
Defining the following
$\textbf{d}_{k}^{(i)}$ $=\begin{bmatrix}\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,\\ \lVert\boldsymbol{\Delta}_{k-1}^{(i)}\rVert\,\end{bmatrix}$ $\textbf{F}_{k}$ $=\begin{bmatrix}\lvert a_{k}\rvert\,\lVert\textbf{H}\rVert\,+\lvert b_{k}\rvert\,+\lvert a_{k}\rvert\,h_{0}&\lvert c_{k}\rvert\,\\ 1&0\end{bmatrix}$ $\textbf{E}_{k}^{(i)}$ $=\begin{bmatrix}\lvert a_{k}\rvert\,h_{1}^{(i)}\\ 0,\end{bmatrix}$
@@
Defining $g_{k}^{0}=\prod_{j=1}^{k}\lVert\textbf{F}_{j}\rVert\,$ we have
$\lVert\boldsymbol{\Delta}_{k+1}^{(i)}\rVert\,$ $\leq g_{k}^{0}\left(\lVert\boldsymbol{\Delta}_{0}^{(i)}\rVert\,+\lVert\boldsymbol{\Delta}_{1}^{(i)}\rVert\,\right)+\sum_{j=1}^{k}\frac{g_{k}^{0}}{g_{j}^{0}}\lvert a_{j}\rvert\,h_{1}^{(i)}$
$=g_{k}^{0}\left(\lVert\boldsymbol{\Delta}_{0}^{(i)}\rVert\,+\lVert\boldsymbol{\Delta}_{1}^{(i)}\rVert\,+\sum_{j=1}^{k}\frac{\lvert a_{j}\rvert\,}{g_{j}^{0}}h_{1}^{(i)}\right)$
Defining $g_{k}^{1}=\sum_{j=1}^{k}\frac{\lvert a_{j}\rvert\,}{g_{j}^{0}}$ we have
$\lVert\boldsymbol{\Delta}_{k+1}^{(i)}\rVert\,$ $\leq g_{k}^{0}\left(\lVert\boldsymbol{\Delta}_{0}^{(i)}\rVert\,+\lVert\boldsymbol{\Delta}_{1}^{(i)}\rVert\,+g_{k}^{1}h_{1}^{(i)}\right)$
Lemma A.4.
@@
where we have defined $f_{k}=\sum_{j=0}^{k}\sum_{r=0}^{j-1}\lVert\alpha_{j}\textbf{H}^{j-r-1}\rVert\,\lVert{\boldsymbol{\Lambda}^{(i)}}^{r}\rVert\,$. ∎
Lemma A.5.
@@
$\lVert\textbf{R}^{(i)}\rVert\,$ $\leq\sin{\angle(\mathcal{S}^{(i)},\mathcal{S})}(\lVert\boldsymbol{\Lambda}_{1}\rVert\,+\lVert\boldsymbol{\Lambda}_{2}\rVert\,)\leq 2\lVert\textbf{H}\rVert\,\sin{\angle(\mathcal{S}^{(i)},\mathcal{S})}$ (18)
B Proof of Theorem 3.2
@@
Using the exact recurrence relation eq. 2 and the recurrence relation employing inexact matrix products eq. 7 we can now write a recurrence relation for $\boldsymbol{\Delta}_{k}^{(i)}=\underline{\textbf{Y}}_{k}^{(i)}-\textbf{Y}_{k}^{(i)}$ as
$\boldsymbol{\Delta}_{k+1}^{(i)}=a_{k}\textbf{H}\boldsymbol{\Delta}_{k}^{(i)}+b_{k}\boldsymbol{\Delta}_{k}^{(i)}+c_{k}\boldsymbol{\Delta}_{k-1}^{(i)}+a_{k}\textbf{e}_{k}^{(i)}$ (19)